Shift Summary for DAY shift, 22/Nov/99

Herman B. White (hwhite@fnal.gov)
Mon, 22 Nov 1999 15:36:51 -0600 (CST)

Shift Summary for Owl shift, 22/Nov/99

On Shift: S.E. Turner, H. B. White

Hours of beam available at all intensities: 8.0
Hours of beam available at nominal intensity: 8.0
Hours of E799 data (special runs not included): 8.0
Hours of special runs (laser scans, muon runs...): 0.0

Accounting of lost time:
We had 26 bad spills one ktevana error.

========================================================================
Runs ended this shift:

Tapes brought to FCC this shift:
none.

========================================================================
Accesses during shift:
none.

========================================================================
Overview of shift:
E799 datataking, Continued run 15030 from last shift without much
interuption. A number of bad spills, see logbook for raw
statistics.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
15030 121 7:18 Split tape request UPF045

246 10:10 misalignment on NM2seed2. beam off corrected
10:22 meson primary beam trip. KteV misaligned
254 problem corrected

15030 12:44 split tape UPH031 request



========================================================================
Follow up Notes on Previous Shift


========================================================================
Items to Follow up on Next Shift
Do a dart_fresh_start for the next run, and a Laser Scan.